A square piece of paper is folded horizontally in half and then cut into 2. If the perimeter of each of the rectangles formed is 36 cm, what is the perimeter of the original square?
1
Expert's answer
2010-07-21T12:29:21-0400
If the perimeter of each of the rectangles is 36 cm we can write the equation: P=(a+b)*2=36 cm
We know that initially a piece of paper was a square, so after folding horizontally in half and then cutting into 2 pieces we have 2 rectangles, in which 1 side is twice larger than the other one, so we can write next equation: P=(a+2a)*2=36 cm; 6a=36 cm; a=6 cm.
The original square piece of paper had side 2a=2*6=12 cm, so the perimeter of the original square is P=a*4=12*4=48 cm.